OpenStudy (anonymous):

Use the quadratic formula to solve the equation. 6x^2 − 7x − 5 = 0 x= (smaller value) x= (larger value)

OpenStudy (anonymous):

\[6x^2-7x-5=0\] \[We find that a=6, b=-7, and c=-5.\] Recall the quadratic formula: \[(-b \pm \sqrt{b^2-4(a)(c)}) /( 2a)\]

OpenStudy (anonymous):

*=x

OpenStudy (anonymous):

?????

OpenStudy (anonymous):

Basically you plug in your numbers. In the equation we can determine A=6, B=-7, and C=-5. Just like the human above pointed out. Then if you look at the quadratic formula which the human above also posted, you just substitute the numbers where necessary. You end up with -(-7)\[\pm\]\[\sqrt{7^{2}-4(6)(-5)}\] and all that divided by 2(6)

OpenStudy (anonymous):

using quadratic formula we get x=-1/2 and x=5/3 ans....